# Implicit Bias of SignGD and Adam on Multiclass Separable Data

### v1

@@ 3 Preliminaries
$\gamma_{\|\cdot\|}:=\max_{\|{\bm{W}}\|\leq 1}\,\min_{\begin{subarray}{c}i\in[n]\\ c\neq{y_{i}}\end{subarray}}\,\left(\bm{e}_{y_{i}}-\bm{e}_{c}\right)^{\top}{\bm{W}}\bm{h}_{i}\,,$ (2)
@@
Optimization Methods
To minimize the empirical loss $\mathcal{L}_{\text{ERM}}$, we use Adam without the stability constant, which performs the following coordinate-wise updates for iteration $t\geq 0$ and initialization ${\bm{W}}_{0}$ (Kingma & Ba, 2014):
${\bm{W}}_{t+1}$ $={\bm{W}}_{t}-\eta_{t}\frac{\mathbf{M}_{t}}{\sqrt{\mathbf{V}_{t}}},$ (3c)
where $\mathbf{M}_{t}$, $\mathbf{V}_{t}$ are first and second moment estimates of the gradient, with decay rates $\beta_{1}$ and $\beta_{2}$. The squaring $(\cdot)^{2}$ and dividing $\frac{\cdot}{\cdot}$ operations are applied entry-wise.
In the special case $\beta_{1}=\beta_{2}=0$, the updates simplify to:
${\bm{W}}_{t+1}$ $={\bm{W}}_{t}-\eta_{t}\texttt{sign}(\nabla\mathcal{L}({\bm{W}}_{t}))),$ (4)
[…]
${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t},\qquad\text{where}$
$\bm{\Delta}_{t}:=\arg\max\nolimits_{\|\bm{\Delta}\|\leq 1}\langle\nabla\mathcal{L}({\bm{W}}_{t}),\bm{\Delta}\rangle\,.$ (5)
Note that this reduces to SignGD, Coordinate Descent (e.g., (Nutini et al., 2015)), or normalized gradient-descent (NGD) when the max-norm (i.e. $p=\infty$), the entry-wise $1$-norm, or the Frobenious Euclidean-norm (i.e. $p=2$) is used, respectively.
Assumptions
Establishing the implicit bias of the above mentioned gradient-based optimization algorithms, requires the following assumptions. First, we assume data are linearly separable, ensuring the margin $\gamma_{\|\cdot\|}$ is strictly positive, an assumption routinely used in previous works (Soudry et al., 2018; Ravi et al., 2024; Soudry et al., 2018; Gunasekar et al., 2018; Nacson et al., 2019; Wu et al., 2024b).
Assumption 3.1.
The second assumption ensures that all entries of the second moment buffer ${\bm{V}}_{t}$ of Adam are bounded away from 00 for all $t\geq 0$. Previously used by Zhang et al. (2024) in binary classification, this assumption is satisfied when the data distribution is continuous and non-degenerate. A similar assumption appears in (Xie & Li, 2024).
Assumption 3.2.
In this work, we consider a decay learning rate schedule of the form $\eta_{t}=\Theta(\frac{1}{t^{a}})$, where $a\in(0,1]$. Such schedule has been studied in the convergence and implicit bias of various optimization algorithms (e.g., (Bottou et al., 2018; Nacson et al., 2019; Sun et al., 2023)) including Adam (Huang et al., 2021; Zhang et al., 2024; Xie & Li, 2024).
Assumption 3.3.
Assumption 3.4 can be satisfied by the above learning rate for a sufficiently large $t$ shown in Zhang et al. (2024, Lemma C.1). It is used in our analysis of Adam.
Assumption 3.4.
Finally, we assume the $1$-norm of the data is bounded. Similar assumptions were used in (Ji & Telgarsky, 2019), (Nacson et al., 2019), (Wu et al., 2024b), and (Zhang et al., 2024).
Assumption 3.5.
4 ${\mathcal{G}}({\bm{W}})$ A proxy to $\mathcal{L}({\bm{W}})$
Analyzing margin convergence begins with studying loss convergence through second-order Taylor expansion of the CE loss:
$\mathcal{L}({\bm{W}}+\bm{\Delta})=\mathcal{L}({\bm{W}})+\langle\nabla\mathcal{L}({\bm{W}}),\bm{\Delta}\rangle$
$\quad+\frac{1}{2n}\sum\nolimits_{i\in[n]}\bm{h}_{i}^{\top}\bm{\Delta}^{\top}{\mathbb{S}}^{\prime}({\bm{W}}\bm{h}_{i})\bm{\Delta}\bm{h}_{i}+o(\|\bm{\Delta}\|_{F}^{3}),$ (6)
where recall that ${\mathbb{S}}^{\prime}({\bm{v}})=\operatorname{diag}({\bm{v}})-{\bm{v}}{\bm{v}}^{\top}$. To bound the loss at ${\bm{W}}_{t+1}={\bm{W}}_{t}-\eta_{t}\bm{\Delta}_{t}$, we must bound both terms in (6). For the NSD updates defined in Eq. (5), the first term evaluates to $-\eta_{t}\|\nabla\mathcal{L}({\bm{W}})\|_{*}$ ($\|\cdot\|_{*}$ is the dual norm). This leads to two key tasks: (1) Lower-bounding the dual gradient norm. (2) Upper-bounding the second-order term.
[…]
Construction of ${\mathcal{G}}({\bm{W}})$
Before showing our construction for the CE loss, it is insightful to discuss how previous works do this in the binary case with labels $y_{b,i}\in{\pm 1}$, classifier vector ${\bm{w}}\in\mathbb{R}^{d}$ and binary margin $\gamma_{b}\coloneqq\max_{\lVert{\bm{w}}\rVert\leq 1}\min_{i\in[n]}y_{b,i}{\bm{w}}^{\top}\bm{h}_{i}$. For exponential loss, Gunasekar et al. (2018) showed that $\lVert\nabla\mathcal{L}({\bm{w}})\rVert\geq\gamma_{b}\mathcal{L}({\bm{w}})$. For logistic loss $\ell(t)=\log(1+\exp(-t))$, Zhang et al. (2024) proved $\lVert\nabla\mathcal{L}({\bm{w}})\rVert_{1}\geq\gamma_{b}{\mathcal{G}}({\bm{w}})$, where ${\mathcal{G}}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$ and $\ell^{\prime}$ is the first-derivative. In both cases, one can take the common form ${\mathcal{G}}_{b}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$. The proof relies on showing $\gamma\leq\min_{\bm{r}\in\triangle^{n-1}}\lVert{\bm{H}}^{T}\bm{r}\rVert$ via Fenchel Duality (Telgarsky, 2013; Gunasekar et al., 2018) and appropriately choosing $\bm{r}$.
In the multiclass setting, where the loss function is vector-valued, it is unclear how to extend the binary proof or definition of ${\mathcal{G}}({\bm{W}})$. To this end, we realize that the key is in the proper manipulation of the gradient inner product $\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle$ (for arbitrary matrix ${\bm{A}}\in\mathbb{R}^{k\times d}$). The CE gradient evaluates to $\nabla\mathcal{L}({\bm{W}})=\frac{1}{n}\sum_{i=1}^{n}(\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i}))\bm{h}_{i}^{\top}$ and using the fact that $\mathbb{S}({\bm{W}}\bm{h}_{i})\in\triangle^{k-1}$, it turns out that we can express (details in Lemma A.1)
$\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle=\frac{1}{n}\sum_{i\in[n]}\sum_{c\neq y_{i}}\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{A}}\bm{h}_{i}\,.$
This motivates defining ${\mathcal{G}}({\bm{W}})$ as:
${\mathcal{G}}({\bm{W}})\coloneqq\frac{1}{n}\sum_{i\in[n]}(1-\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i}))\,.$ (7)
The following key lemma , which directly follows from the inner-product calculation above and our definition of ${\mathcal{G}}({\bm{W}})$, confirms that this is the right choice. For convenience, denote $s_{ic}\coloneqq\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})$, for $i\in[n],c\in[k]$.
Lemma 4.1 (Lower bounding the gradient dual-norm).
For any ${\bm{W}}\in\mathbb{R}^{k\times d}$, it holds that ${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})$.
Proof.
By duality, the above calculated formula for the CE-gradient inner product and the fact that $\sum_{c\in[k]}s_{ic}=1$:
${\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}=\max\nolimits_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\,\langle\bm{A},-\nabla\mathcal{L}({\bm{W}})\rangle$
$=\max\nolimits_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\,\frac{1}{n}\sum\nolimits_{i\in[n]}\sum\nolimits_{c\neq y_{i}}s_{ic}\,(\bm{e}_{{y_{i}}}-\bm{e}_{c})^{\top}\bm{A}\bm{h}_{i}$
$\geq\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\,\min_{c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}$
$=\frac{1}{n}\sum_{i\in[n]}(1-s_{iy_{i}})\cdot\max_{{\left\|\bm{A}\right\|_{\max}}\leq 1}\min_{i\in[n],c\neq y_{i}}(\bm{e}_{y_{i}}-\bm{e}_{c})^{T}\bm{A}\bm{h}_{i}\,.$
To finish recall the definitions of ${\mathcal{G}}({\bm{W}})$ and of the max-norm margin $\gamma$. ∎
The lemma completes the first of our two tasks: lower bounding the gradient’s dual norm. Importantly, the factor appearing in front of ${\mathcal{G}}({\bm{W}})$ is the max-norm margin $\gamma$, which will prove crucial in the forthcoming margin analysis.
${\mathcal{G}}({\bm{W}})$ and second-order term
We now show how to bound the second-order term in (6). For this, we establish the following essential lemma.
Lemma 4.2.
For any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and any ${\bm{v}}\in\mathbb{R}^{k}$ it holds:
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4\,(1-s_{c})\,\|{\bm{v}}\|_{\infty}^{2}\,.$
Proof.
By Cauchy-Schwartz,
${\bm{v}}^{\top}$ $\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}=\langle\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top},{\bm{v}}{\bm{v}}^{\top}\rangle$
$\leq\,{\left\|\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right\|_{\rm{sum}}}\,\|{\bm{v}}\|_{\infty}^{2}\,.$
Direct calculation yields ${\left\|\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right\|_{\rm{sum}}}=2\sum\nolimits_{c\in[k]}s_{c}(1-s_{c})$. The advertised bound then follows by noting the following $\sum_{c\in[k]}s_{c}(1-s_{c})\leq 2(1-s_{c^{\prime}})$ for any $c^{\prime}\in[k]$ (verified in Lemma A.4). ∎
To bound the second-order term in Eq. (6), we can apply the above lemma with ${\bm{v}}\leftarrow\bm{\Delta}\bm{h}_{i}$ and $c\leftarrow y_{i}$ and further use the inequality $\|{\bm{\Delta}\bm{h}_{i}}\|_{\infty}\leq{\left\|\bm{\Delta}\right\|_{\max}}\|\bm{h}_{i}\|_{1}\leq B{\left\|\bm{\Delta}\right\|_{\max}}$ (the last step invoked Ass. 3.5). This yields an upper bound
$2B^{2}\|{\bm{\Delta}}\|_{\max}^{2}\cdot\frac{1}{n}\sum_{i\in[n]}(1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i}))$
for the second-order term in the CE loss expansion in terms of the proxy function ${\mathcal{G}}({\bm{W}})$. Thus, we have fullfilled the first desiderate by finding ${\mathcal{G}}({\bm{W}})$ that simultaneously bounds the first and second-order terms in Eq. (6).
@@ 4 ${\mathcal{G}}({\bm{W}})$ A proxy to $\mathcal{L}({\bm{W}})$
We now show that the function ${\mathcal{G}}({\bm{W}})$ in Eq. (7) meets the second desiderata: being a good proxy for the loss $\mathcal{L}({\bm{W}})$. This is rooted in elementary relationships between ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$, which play crucial roles in the various parts of the proof. Below, we summarize this key relationships.
Lemma 4.3 (Properties of ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$. The following hold: (i) Under Ass. 3.5, $2B\cdot{\mathcal{G}}({\bm{W}})\geq{\left\|\nabla\mathcal{L}({\bm{W}})\right\|_{\rm{sum}}}$; (ii) $1\geq\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\geq 1-\frac{n\mathcal{L}({\bm{W}})}{2}$; (iii) If ${\bm{W}}$ satisfies $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ or ${\mathcal{G}}({\bm{W}})\leq\frac{1}{2n}$, then $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}})$.
Lemma 4.3 (i) extends Lemma 4.1 by establishing a sandwich relationship between ${\mathcal{G}}({\bm{W}})$ and the gradient’s dual norm. The lemma’s statements (ii) and (iii) show that ${\mathcal{G}}({\bm{W}})$ can substitute for the loss - it lower bounds $\mathcal{L}({\bm{W}})$ and serves as an upper bound when either $\mathcal{L}({\bm{W}})$ or ${\mathcal{G}}({\bm{W}})$ is sufficiently small. Specifically, the ratio ${{\mathcal{G}}({\bm{W}})}\big{/}{\mathcal{L}({\bm{W}})}$ converges to 1 as the loss decreases, with the convergence rate depending on the rate of decrease. The key property (ii) may seem algebraically complex, but it turns out (Lemma B.3) that both sides of the sandwich relationship follow from the elementary fact that $\forall x>0:1-x\leq e^{-x}\leq 1-x+x^{2}/2$.
5 Implicit Bias of SignGD
We now leverage our construction of ${\mathcal{G}}({\bm{W}})$ to show that the margin of SignGD’s iterates converges to max-norm margin. We only highlight the key steps in the proof and defer details to Appendix C.
SignGD Descent
We start by showing a descent property. By applying Lemmas 4.1 and 4.2 to lower and upper bound the first and second order terms in Eq. (6) yields:
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+$
$\quad\quad 2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t})\sup_{\zeta\in[0,1]}\frac{{\mathcal{G}}({\bm{W}}_{t}+\zeta\bm{\Delta}_{t})}{{\mathcal{G}}({\bm{W}}_{t})}.$
Algebraic manipulations of the definition of ${\mathcal{G}}({\bm{W}})$ allows us to bound the ratio in the right hand side.
Lemma 5.1 (Ratio of ${\mathcal{G}}({\bm{W}})$).
For any $\psi\in[0,1]$, we have the following: $\frac{{\mathcal{G}}({\bm{W}}+\psi\triangle{\bm{W}})}{{\mathcal{G}}({\bm{W}})}\leq e^{2B\psi{\left\|\triangle{\bm{W}}\right\|_{\max}}}$.
From this and the fact ${\left\|\bm{\Delta}_{t}\right\|_{\max}}\leq\eta_{t}$ for SignGD, we obtain

### v4

@@ 3 A Unified Framework with a Proxy Function
These relationships (proved in Lemma 11 in App. C) are crucial for unifying the analysis of NSD and NMD algorithms w.r.t. either the entry-wise or the Schatten norms (details below).
Construction of ${\mathcal{G}}({\bm{W}})$
Before showing our construction for the CE loss, it is insightful to discuss how previous works do this in the binary case with labels $y_{b,i}\in\{\pm 1\}$, classifier vector ${\bm{w}}\in\mathbb{R}^{d}$ and binary margin $\gamma_{b}\coloneqq\max_{\lVert{\bm{w}}\rVert\leq 1}\min_{i\in[n]}y_{b,i}{\bm{w}}^{\top}\bm{h}_{i}$. For exponential loss, gunasekar2018characterizing showed that $\lVert\nabla\mathcal{L}({\bm{w}})\rVert\geq\gamma_{b}\mathcal{L}({\bm{w}})$. For logistic loss $\ell(t)=\log(1+\exp(-t))$, zhang2024implicit proved $\lVert\nabla\mathcal{L}({\bm{w}})\rVert_{1}\geq\gamma_{b}{\mathcal{G}}({\bm{w}})$, where ${\mathcal{G}}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$ and $\ell^{\prime}$ is the first-order derivative. In both cases, one can take the common form ${\mathcal{G}}_{b}({\bm{w}})=\frac{1}{n}\sum_{i=1}^{n}|\ell^{\prime}(y_{b,i}{\bm{w}}^{\top}\bm{h}_{i})|$. The proof relies on showing $\gamma\leq\min_{\bm{r}\in\triangle^{n-1}}\lVert{\bm{H}}^{T}\bm{r}\rVert$ via Fenchel Duality [telgarsky2013margins, gunasekar2018characterizing] and appropriately choosing $\bm{r}$.
In the multiclass setting, where the loss function is vector-valued, it is unclear how to extend the binary proof or definition of ${\mathcal{G}}({\bm{W}})$. To this end, we realize that the key is in the proper manipulation of the gradient inner product $\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle$ (for arbitrary matrix ${\bm{A}}\in\mathbb{R}^{k\times d}$). The CE gradient evaluates to $\nabla\mathcal{L}({\bm{W}})=\frac{1}{n}\sum_{i=1}^{n}(\bm{e}_{y_{i}}-\mathbb{S}({\bm{W}}\bm{h}_{i}))\bm{h}_{i}^{\top}$ and using the fact that $\mathbb{S}({\bm{W}}\bm{h}_{i})\in\triangle^{k-1}$, it turns out that we can express (details in Lemma 9): $\langle{\bm{A}},-\nabla\mathcal{L}({\bm{W}})\rangle=\frac{1}{n}\sum\nolimits_{i\in[n]}\sum\nolimits_{c\neq y_{i}}\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})(\bm{e}_{y_{i}}-\bm{e}_{c})^{\top}{\bm{A}}\bm{h}_{i}\,.$
This motivates defining ${\mathcal{G}}({\bm{W}})$ as:
${\mathcal{G}}({\bm{W}})\coloneqq\frac{1}{n}\sum\nolimits_{i\in[n]}(1-\mathbb{S}_{{y_{i}}}({\bm{W}}\bm{h}_{i}))\,.$ (8)
The lemma below, following from the inner-product calculation above and our definition of ${\mathcal{G}}({\bm{W}})$, confirms this is the right choice. For convenience, denote $s_{ic}\coloneqq\mathbb{S}_{c}({\bm{W}}\bm{h}_{i})$, for $i\in[n],c\in[k]$.
Lemma 1 (Lower bounding the gradient dual-norm).
For any ${\bm{W}}\in\mathbb{R}^{k\times d}$ and any entry-wise or Schatten p-norm $\lVert\cdot\rVert$ with $p\geq 1$, it holds that $\lVert\nabla\mathcal{L}({\bm{W}})\rVert_{*}\geq\gamma\cdot{\mathcal{G}}({\bm{W}})$, where $\lVert\cdot\rVert_{*}$ is the dual-norm.
The lemma completes the first task: lower bounding the gradient’s dual norm. Importantly, the factor in front of ${\mathcal{G}}({\bm{W}})$ is the margin $\gamma$ w.r.t. the norm $\lVert\cdot\rVert$, which is crucial in the forthcoming analysis.
${\mathcal{G}}({\bm{W}})$ and second-order term
We now show how to bound the second-order term in (6). For this, we establish the following essential lemma whose proof relies on the relationships in (7).
Lemma 2.
For any entry-wise or Schatten p-norm $\lVert\cdot\rVert$ with $p\geq 1$, any $\bm{s}\in\Delta^{k-1}$ in the $k$-dimensional simplex, any index $c\in[k]$, and ${\bm{v}}\in\mathbb{R}^{k}$, it holds that
${\bm{v}}^{\top}\left(\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}\right){\bm{v}}\leq 4(1-s_{c})\lVert{\bm{v}}{\bm{v}}^{T}\rVert.$
Proof.
Let $\bm{S}:=\operatorname{diag}(\bm{s})-\bm{s}\bm{s}^{\top}$ and $q\geq 1$ such that $1/p+1/q=1$. By norm duality, it holds that
${\bm{v}}^{\top}\bm{S}{\bm{v}}=\operatorname{tr}\left({\bm{S}{\bm{v}}{\bm{v}}^{\top}}\right)\leq\|\bm{S}\|_{q}\|{\bm{v}}{\bm{v}}^{\top}\|\leq{\left\|\bm{S}\right\|_{\rm{sum}}}\|{\bm{v}}{\bm{v}}^{\top}\|,$
where $\lVert\cdot\rVert_{q}$ is the dual of $\lVert\cdot\rVert$ and the second inequality is by (7). Direct calculation yields ${\left\|\bm{S}\right\|_{\rm{sum}}}=2\sum\nolimits_{c\in[k]}s_{c}(1-s_{c})$. The advertised bound then follows by noting the following $\sum_{c\in[k]}s_{c}(1-s_{c})\leq 2(1-s_{c^{\prime}})$ for any $c^{\prime}\in[k]$ (verified in Lemma 13 in App. C). ∎
@@
$2B^{2}\|{\bm{\Delta}}\|^{2}\cdot\frac{1}{n}\sum\nolimits_{i\in[n]}(1-\mathbb{S}_{y_{i}}({\bm{W}}\bm{h}_{i}))\,.$
@@
We now show that ${\mathcal{G}}({\bm{W}})$ meets the second desiderata: being a good proxy for the loss $\mathcal{L}({\bm{W}})$. This is rooted in the elementary relationships between ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$, which are used in the various parts of the proof. Below, we summarize these key relationships.
Lemma 3 (Properties of ${\mathcal{G}}({\bm{W}})$ and $\mathcal{L}({\bm{W}})$).
Let ${\bm{W}}\in\mathbb{R}^{k\times d}$. The followings hold: (i) Under Ass. 4, $2B\cdot{\mathcal{G}}({\bm{W}})\geq\lVert\nabla\mathcal{L}({\bm{W}})\rVert_{*}$; (ii) $1\geq\frac{{\mathcal{G}}({\bm{W}})}{\mathcal{L}({\bm{W}})}\geq 1-\frac{n\mathcal{L}({\bm{W}})}{2}$; (iii) If $\bm{W}$} satisfies $\mathcal{L}({\bm{W}})\leq\frac{\log 2}{n}$ or ${\mathcal{G}}({\bm{W}})\leq\frac{1}{2n}$, then $\mathcal{L}({\bm{W}})\leq 2{\mathcal{G}}({\bm{W}})$.
Lemma 3 (i) extends Lemma 1 by establishing a sandwich relationship between ${\mathcal{G}}({\bm{W}})$ and the gradient’s dual norm. The lemma’s statements (ii) and (iii) show that ${\mathcal{G}}({\bm{W}})$ can substitute for the loss - it lower bounds $\mathcal{L}({\bm{W}})$ and serves as an upper bound when either $\mathcal{L}({\bm{W}})$ or ${\mathcal{G}}({\bm{W}})$ is sufficiently small. Specifically, the ratio ${{\mathcal{G}}({\bm{W}})}\big/{\mathcal{L}({\bm{W}})}$ converges to 1 as the loss decreases, with the convergence rate depending on the rate of loss decrease. The key property (ii) may seem algebraically complex, but it turns out (details in Lemma 18 in App. C) that both sides of the sandwich relationship follow from the elementary fact that $\forall x>0:1-x\leq e^{-x}\leq 1-x+x^{2}/2$.
4 Implicit Bias of Normalized Steepest Descent
We now leverage our construction of ${\mathcal{G}}({\bm{W}})$ to show that the margin of NSD’s iterates converges to the data margin defined w.r.t. the same entry-wise or Schatten p-norm that is used to define the algorithm (refer to eqns. (2) and (3) for the definitions of margin and NSD). We only highlight the key steps in the proof and defer details to App. D.
NSD Descent
We start by showing a descent property. By applying Lemmas 1 and 2 to lower and upper bound the first and second order terms in Eq. (6) yields
$\mathcal{L}({\bm{W}}_{t+1})$ $\leq\mathcal{L}({\bm{W}}_{t})-\gamma\eta_{t}{\mathcal{G}}({\bm{W}}_{t})+2\eta_{t}^{2}B^{2}{\mathcal{G}}({\bm{W}}_{t})\sup_{\zeta\in[0,1]}\frac{{\mathcal{G}}({\bm{W}}_{t}-\zeta\eta_{t}\bm{\Delta}_{t})}{{\mathcal{G}}({\bm{W}}_{t})}.$
Algebraic manipulations of the definition of ${\mathcal{G}}({\bm{W}})$ and the relationships in (7) allow us to bound the ratio in the right hand side.
Lemma 4 (Ratio of ${\mathcal{G}}({\bm{W}})$).
For any $\psi\in[0,1]$, we have the following: $\frac{{\mathcal{G}}({\bm{W}}-\psi\eta\bm{\Delta})}{{\mathcal{G}}({\bm{W}})}\leq e^{2B\eta\psi{\left\|\bm{\Delta}\right\|_{\max}}}\leq e^{2B\eta\psi\lVert\bm{\Delta}\rVert}$ (note that the second inequality is by (7)).
From this and $\lVert\bm{\Delta}_{t}\rVert\leq 1$ for NSD, we obtain